A particle is moving eastward with a velocity of 5m/s . In 10 sec the velocity changes to 5m/s nortward . The average acceleration delta v/delta t in this time is:
a) zero                      
b)1/root 2 m/s square towards north west
c)1/root 2 m/s square towards north east
1/root 2 m/s square towards north

Grade:11

1 Answers

Arun
4 years ago
Intial velocity=v1= 5m/seast wards                                   ^||^
then after time t=10sec                                                     || towards north 5m/s
v2 =5m/s north wards                                     .⇒⇒⇒⇒⇒⇒|
                                                                towards east 5m/s
Vresultant = √v1²+v2²
                =√25+25
                 =5 √2
acceleration a=velocity/time
                     = 5√2/10 
                     =√2/2 × √2/√2
                     =1/√2 towards n- w